Interim Constitution of the SHIP Collaboration – Appendix B

(7 October 2015) Member Institutes

1. University of Sofia

2. UTFSM (Universidad Tcnica Federico Santa), Valparaiso

3. NBI (Niels Bohr Institute), CopenHagen University

4. LAL Orsay

5. LPNHE Univ. Paris 6 et 7

6. Humboldt University of Berlin

7. University of Hamburg

8. Istituto Nazionale di Fisica Nucleare (INFN), Sezione di Bari

9. Istituto Nazionale di Fisica Nucleare (INFN), Sezione di Bologna

10. Istituto Nazionale di Fisica Nucleare (INFN), Sezione di Cagliari

11. Istituto Nazionale di Fisica Nucleare (INFN), Sezione di Ferrara

12. Lab. Naz. Gran Sasso

13. Lab. Naz. Frascati

14. Istituto Nazionale di Fisica Nucleare (INFN), Sezione di Napoli

15. University of Rome La Sapienza

16. Aichi University of Education

17. Kobe University

18. Nagoya University

19. Nihon University

20. Toho University

21. Gyeongsang National University

22. Joint Institute of Nuclear Research (JINR) Dubna

23. Institute for Theoretical and Experimental Physics (ITEP), Moscow

24. Institute for Nuclear Research (INR), Moscow

25. P.N. Lebedev Physical Institute of the Russian Academy of Sciences (LPI), Moscow 26. National Research Centre (NRC) ”Kurchatove Institute”, Moscow

27. Institute for High Energy Physics, Protvino

28. Petersburg Nuclear Physics Institute (PNPI), St. Petersburg 29. Moscow Engineering Physics Institute (MEPhI)

30. Skobeltsyn Institute of Nuclear Physics of Moscow State University

31. Yandex School of Data Analysis

32. Stockholm University

33. Uppsala University

34. CERN

35. University of Geneva

36. Ecole Polytechnique Federale de Lausanne (EPFL)

37. University of Zürich

38. Middle East Technical University (METU), Ankara

39. Ankara University

40. Imperial College London

41. University College London

42. Rutherford Appleton Laboratory (RAL)

43. Bristol University

44. Warwick University

45. Taras Shevchenko National University of Kyiv

46. Florida University Associated Member Institutes

1. National University of Science and Technology ”MISIS”, Moscow. Associated to LPI, Moscow.